OpenStudy (cggurumanjunath):

prove\quad that\quad 2{ n }^{ 2 }\ge { \left( n+1 \right) }^{ 2 }\quad for\quad n\ge 3

OpenStudy (cggurumanjunath):

\[prove\quad that\quad 2{ n }^{ 2 }\ge { \left( n+1 \right) }^{ 2 }\quad for\quad n\ge 3\]

OpenStudy (cggurumanjunath):

can anybody help me with this problem of mathematical analysis.

myininaya (myininaya):

So you try this by induction?

OpenStudy (cggurumanjunath):

s i don't know how to proceed , what i did was i just substituted n=3 on both sides and found that 18>16.i don't know how to proceed further.

OpenStudy (anonymous):

For any value n= k lets assume 2k^2>=(k+1)^2 then For n= k+1 , 2(k+1)^2>=(k+1+1)^2 2(k^2+2k+1) >= (k+1)^1 + 2(k+1) +1 2k^2 +4k +1 >= (k+1)^2 +2k +3 2k^2 +2k >= (k+1)^2 + 2 Now, when k=3 , 2*3^2 >= (3+1)^2 =>18 >=16 [which is true] and 2k >= 2 is also true when k>=3 So, 2k^2+2k >= (k+1)^2 +2 is also true for all k>=3 Thus, 2n^2 >= (n+1)^2 for n>=3

OpenStudy (cggurumanjunath):

thank you very much @sauravshakya.